OpenStudy (calculusxy):

Medal! A cookie recipe requires 3/4 cup of flour and 1/3 cup of butter to make one batch of cookies. Wendy has 2 cups of flour and 3/4 cup of butter. How many whole batches of cookies can she make? A. 1 B. 2 C. 3 D. 4

OpenStudy (anonymous):

wait while I do an equation.

OpenStudy (anonymous):

Flour: \[\frac{ 3 }{ 4 }+\frac{ 3 }{ 4 }=\frac{ 6 }{ 4 }\]Flour totals to 1 1/2 cups Butter: \[\frac{ 1 }{ 3 }+\frac{ 1 }{ 3 }=\frac{ 2 }{ 3 }\] Butter totals to 2/3 cups. Can you figure out how many batches she can make?

OpenStudy (calculusxy):

Do you add the 2/3 and 6/4?

OpenStudy (anonymous):

Hint: 3/4 is bigger than 2/3

OpenStudy (anonymous):

No, think about the numbers separately.

OpenStudy (anonymous):

How many batches of flour can she have with two cups of flour.

OpenStudy (anonymous):

How many batches of butter can she have with 3/4 cups of butter.

OpenStudy (calculusxy):

I was thinking about dividing the 2 by 3/4 (flour) and 3/4 divided by 1/3 (butter). Then add the whole numbers that there are.

OpenStudy (anonymous):

She can make two batches of cookies, because she has enough of each ingredient for both batches.

OpenStudy (calculusxy):

Do you figure that out by dividing? Or by your strategy?

OpenStudy (calculusxy):

When I divided 2 by 3/4 = 2 2/3 When I divided 3/4 by 1/3 = 2 1/4 Add the whole numbers = 2 + 2 = 4

OpenStudy (anonymous):

I just used my own strategy. I added them together to figure out how much would be needed for two recipes.